# Agent Journey Beyond RGB: Unveiling Hybrid Semantic-Spatial Environmental Representations for Vision-and-Language Navigation

### v2

@@ 3.4 Hybrid Representation Fusion and Prediction
^$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\mathcal{X}^{\mathcal{S}},\mathcal{T}^{\mathcal{D}},\mathcal{X}^{R},\mathcal{T}^{R},\mathcal{I}\},$ (4)
@@
with $\tau$ representing the tanh activation function. Subsequently, we perform weighted fusion of the hybrid environmental representations $E^{\mathrm{hyb}}$, as follows:
$\small\footnotesize\mathcal{B}=\begin{bmatrix}\beta_{1}&\beta_{2}&\beta_{3}&\beta_{4}\end{bmatrix}^{\top}=\sigma(\mathrm{FFN}[\mathcal{X}_{0}^{\mathcal{S}};\mathcal{X}_{0}^{\mathcal{R}};\mathcal{T}_{0}^{\mathcal{D}};\mathcal{T}_{0}^{\mathcal{R}}])$ (6)
$\small\footnotesize E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B}$ (7)
@@ 3.5 Training and Inference
Pretraining. Our pre-training phase followed prior work [5, 25], utilizing auxiliary tasks such as Masked Language Modeling (MLM), Masked Region Classification (MRC), and Single-step Action Prediction (SAP). Object Grounding (OG) task to localize target objects in REVERIE and SOON. Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-train the model may lead to overfitting and increased training costs. As shown in Fig. 2, only the computation flows indicated by black arrows were pre-trained. The effectiveness of this partial pre-trained strategy is demonstrated in Sec. 4.3.
@@
$\textstyle\mathcal{L}_{cl}=-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}{\sum_{j=1}^{B}\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}$ (10)
$\textstyle-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{j})/t)}{\sum_{j=1}^{B}\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{j})/t)},$
where $B$ and $t$ represent the batch size and the temperature. Then, we train the model based on hybrid environment representations by performing single-step action prediction [4, 5] with a contrastive learning loss $\mathcal{L}_{cl}$. The hybrid single action prediction loss in behavior cloning is computed as $\mathcal{L}_{hsap}^{*}=\sum_{t=1}^{T}-\log p(a_{t}^{*}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{*}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map. The final training loss is then expressed as follows:
$\mathcal{L}_{SUSA}=\lambda_{1}\mathcal{L}_{hsap}^{*}+\mathcal{L}_{hsap}^{\star}+\lambda_{2}\mathcal{L}_{cl}$ (11)
@@
During the inference phase, the agent is tasked with sequentially navigating to the predicted navigable points, following the shortest path algorithm [12], without making any abrupt “jumps” between points.
@@ 4.1 Task Setup and Implementation
Metrics. We evaluate the navigation performance using standard metrics: Trajectory Length (TL), Success Rate (SR), Navigation Error (NE), and Success Rate weighted by Path Length (SPL). Additionally, we use Normalized Dynamic Time Warping (nDTW) and Success weighted by Normalized Dynamic Time Warping (sDTW) to evaluate the fidelity of successful trajectories to their corresponding instructions. For the REVERIE and SOON tasks, which involve object identification, we also evaluate object grounding metrics: Remote Grounding Success (RGS), and RGS weighted by Path Length (RGSPL).
Implementation Details. We initialized transformer layers with pretrained LXMERT [49] during pre-training. Pre-training was performed for 100k iterations with a batch size of 32. Fine-tuning was performed for 25k iterations. For fine-tuning, we used batch sizes of 4, 8, and 2 for R2R, REVERIE, and SOON, respectively. The default hyperparameters are $\lambda_{1}$ = 0.2 and $\lambda_{2}$= 0.8. The node and view dimensions are identical: $d_{v}=d_{n}=d=768$. All experiments were conducted on a single NVIDIA RTX 4090 GPU. We minimized changes to the baseline [5] settings and used no additional annotations. Detailed experimental setup is provided in the supplementary material.
@@ 4.2 Comparisons with State of the Art
We conduct a comprehensive comparison on the REVERIE [40], R2R [3], and SOON [65] datasets. For fairness, we exclude methods that involve pre-exploration [66, 46] or utilize large-scale data augmentation [6]. REVERIE. Table 1 presents a comparative analysis on the REVERIE. On the test unseen split, SUSA achieves an SPL score of 41.54%, significantly outperforming the previous state-of-the-art (SOTA) method by a large margin of 2.78%. Furthermore, we establish new SOTA results for RGS and RGSPL on both the validation and test unseen splits, demonstrating that the introduced textual semantics substantially elevate overall navigation, particularly benefiting object localization performance. R2R. The compared experiment results on R2R, summarized in  Tab. 2. Compared with the SOTA model GOAT [54] on the validation and test unseen splits, SUSA achieves 64.85% and 63.83% on SPL, respectively. These gains manifest not only an improved success rate but also enhanced navigation efficiency, since SPL considers both success rate and trajectory length.
@@
SOON. On the SOON dataset, as evidenced in  Tab. 3, SUSA outperforms the SOTA model GOAT by 2.27% and 0.29% on the validation and test unseen splits, respectively, in terms of SPL. Overall, the outstanding experimental results show that textual semantics boost navigation success via salient landmark matching, while the depth exploration map is beneficial to generalizable navigation efficiency.
@@ 4.3 Diagnostic Experiment
1 13.41 3.10 79.86 71.43 60.63
2 13.85 3.08 80.42 72.80 62.71
3 13.78 3.06 81.65 72.92 63.16
4 12.18 3.06 79.99 73.05 64.85
5 14.58 3.08 82.08 73.95 62.84
To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the R2R validation unseen split in the Tab. 4. As shown in (#2, #3, #4), both the Textual-aware Semantic Understanding (TSU) and Depth-aware Spatial Perception (DSP) modules consistently outperform the baseline (#1) [5], underscoring the substantial performance gains achieved by incorporating textual panoramas and depth exploration maps over relying solely on RGB. The hybrid representation fusion (HRF) module improves the OSR/SR score from 73.05%/79.99% to 73.95%/82.08%, demonstrating that the proposed HRF enables more precise alignment of environment-instruction pairs through contrastive learning. Analysis of the HRF module and the computational efficiency of the entire model is presented in supplementary materials.
[…]
ViT No 51.07 46.98 33.73 32.15 23.03
CLIP No 57.17 51.55 35.75 35.19 24.40
CLIP ResNet(ImageNet) 54.42 50.18 34.63 34.39 23.66
CLIP ResNet(Gibson) 56.72 52.00 39.21 35.05 26.54
Table 5: Ablation of different visual representations. Gibson and ImageNet serve as pre-training sources for the depth encoder.
@@
2) Image Features vs. Structural Features. We analyze various visual features (depth and RGB) contributions within the DSP module (Tab. 5). For RGB images, substituting ViT [10] features with those from CLIP [43] yields marked improvements across all metrics, underscoring that CLIP achieves a better understanding of the environment than its ViT counterpart. As for depth features, we leveraged depth features extracted by ResNet-50 [16] pre-trained on the ImageNet [8] or Gibson [59] datasets to investigated the effect of depth exploration map. Compared to Matterport3D, the Gibson simulator offers more detailed and realistic physical environments along with higher-resolution visual information. The results demonstrate that simply incorporating depth information does not necessarily lead to improved performance. In comparison, the ResNet-50 pre-trained on the Gibson dataset extracts structural features that refine spatial perception, yielding notable improvements in the SPL metric without compromising SR or RGS. These improvements indicate that better spatial perception enhances navigation efficiency.
@@
3) Static Matching vs. Dynamic Matching. To assess the impact of static matching ($\delta=0$) and dynamic matching ($\delta=1$) strategies in the TSU module, we perform ablation studies based on balance factor $\delta$ on the R2R and REVERIE.  Tab. 6 summarizes these results, where w/o indicates the exclusion of the TSU module, corresponding to the results in Tab. 4 (#2) and last row in Tab. 5. Furthermore, adaptive denotes that $\delta$ is a learnable parameter which adaptively adjusts the contribution of each matching strategy. Compared to the first row in Tab. 6, overall, incorporating the TSU module (regardless of static or dynamic matching strategy) improves SR and RGS, albeit at the cost of SPL on REVERIE. In contrast, it enhances SPL on R2R. While the TSU module can achieve superior performance under the adaptive setting, it falls short when $\delta$ = 0.5. Specifically, at $\delta$ = 0.5, the SPL metric reaches 39.50% and 64.85% on the two datasets, with the RGS, nDTW, and sDTW metrics also achieving optimal performance. Excellent results demonstrate dual merits of proposed TSU module, which boosts object identification accuracy on the REVERIE task while simultaneously improving the fidelity between navigation trajectories and instructions on the R2R.
@@
Figure 5: a) and b) show the performance gap ($\downarrow$) between seen and unseen environments, while c) and d) present key metrics for different pretraining strategies on the R2R and REVERIE.
@@
4) Improved Generalization via Pretraining. As analyzed before, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which can negatively impact the generalization in unseen environments, particularly considering the modality differences between textual panorama, depth exploration maps, and RGB representations. As shown in Fig. 5, we observe that partial pretraining strategies not only consistently narrow the performance gap between seen and unseen environments but also outperform the completely/random pretraining strategies in navigation performance. For example, as depicted in Fig. 5 (a), on the REVERIE, the partial pretraining model achieves an SR gap of only 21.11% between the seen and unseen splits, reducing the gap by 5.92% compared to the complete pretraining strategy. Additionally, as shown in Fig. 5 (c), multiple metrics indicate that partial pretraining results in superior performance. Fig. 5(b) and (d) illustrate similar trends are observed on R2R, further emphasizing the impressive generalization power of the partial pretraining strategy.
@@ 4.4 Qualitative Analysis
Figure 6: Predicted trajectories of SUSA and DEUT on the R2R validation unseen split.
@@ 5 Conclusion
We present the SUSA architecture, which offers complementary environmental representations beyond RGB images. We propose the TSU module, enabling the agent to identify the most instruction-relevant views through both static and dynamic matching. The DSP module enhances spatial awareness by integrating depth exploration maps. We further leverage contrastive learning to precisely align hybrid environmental representations with instructions. Experimental results on the R2R, REVERIE, and SOON benchmarks demonstrate the effectiveness and generalization of our approach. Future work will enrich beneficial environmental representations to advance specific VLN tasks.
@@ 6 Detailed Experimental Setups
We evaluated the experimental results of the unseen test split on the public online leaderboards***https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate. We assure the readers that we will make our code and model checkpoints publicly available at https://github.com/HCI-LMC/VLN-SUSA.

### v4

@@ 3.4 Hybrid Representation Aggregation
^$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\tilde{\mathcal{X}^{\mathcal{S}}},\tilde{\mathcal{T}^{\mathcal{D}}},\tilde{\mathcal{X}^{R}},\tilde{\mathcal{T}^{R}},\mathcal{I}\},$ (4)
@@
with $\tau$ representing the tanh activation function. Subsequently, we perform weighted aggregation of the hybrid environmental representations $E^{\mathrm{hyb}}$, as follows:
$\small\footnotesize\mathcal{B}=\begin{bmatrix}\beta_{1}&\beta_{2}&\beta_{3}&\beta_{4}\end{bmatrix}^{\top}=\sigma(\mathrm{FFN}[\tilde{\mathcal{X}_{0}^{\mathcal{S}}};\tilde{\mathcal{X}_{0}^{\mathcal{R}}};\tilde{\mathcal{T}_{0}^{\mathcal{D}}};\tilde{\mathcal{T}_{0}^{\mathcal{R}}}])$ (6)
$\small\footnotesize E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B},$ (7)
@@ 3.5 Training and Inference
Pretraining. Our pre-training phase followed prior work [6, 30], utilizing auxiliary tasks such as Masked Language Modeling (MLM), Masked Region Classification (MRC), and Single-step Action Prediction (SAP). Object Grounding (OG) task to localize target objects in REVERIE and SOON. Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-train all parameter of SUSA may lead to overfitting and increased training costs. As shown in Fig. 2, only the computation flows indicated by black arrows were pre-trained. The effectiveness of this partial pre-trained strategy is demonstrated in Sec. 4.3.
@@
$\textstyle\mathcal{L}_{cl}=-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{i})/t)}{\sum_{j=1}^{B}\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}$ (10)
$\textstyle-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{i})/t)}{\sum_{j=1}^{B}\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{j})/t)},$
where $B$ and $t$ represent the batch size and the temperature. Then, we train the model based on hybrid environment representations by performing single-step action prediction [5, 6] with a contrastive learning loss $\mathcal{L}_{cl}$. The hybrid single action prediction loss in behavior cloning is computed as $\mathcal{L}_{hsap}^{gt}=\sum_{t=1}^{T}-\log p(a_{t}^{gt}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{gt}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map. The final training loss is then expressed as follows:
$\mathcal{L}_{SUSA}=\lambda_{1}\mathcal{L}_{hsap}^{gt}+\mathcal{L}_{hsap}^{\star}+\lambda_{2}\mathcal{L}_{cl}$ (11)
@@
During the inference phase, the agent is tasked with sequentially navigating to the predicted navigable nodes, following the shortest path algorithm [13], without making any abrupt “jumps” between nodes.
@@ 4.1 Task Setup and Implementation
Metrics. We evaluate the navigation performance using standard metrics: Trajectory Length (TL), Success Rate (SR), Navigation Error (NE), and Success Rate weighted by Path Length (SPL). Additionally, we use Normalized Dynamic Time Warping (nDTW) and Success weighted by Normalized Dynamic Time Warping (sDTW) to evaluate the fidelity of successful trajectories to their corresponding instructions. For the REVERIE and SOON tasks, which involve object identification, we also evaluate object grounding metrics: Remote Grounding Success (RGS), and RGS weighted by Path Length (RGSPL). Given that SPL optimally balances navigation success rate and trajectory length, we employ it as our primary metric.
Implementation Details. We initialized transformer layers with pretrained LXMERT [48] during pre-training. Pre-training was performed for 100k iterations with a batch size of 32. Fine-tuning was performed for 25k iterations. For fine-tuning, we used batch sizes of 4, 8, and 2 for R2R, REVERIE, and SOON, respectively. The default hyperparameters are $\lambda_{1}$ = 0.2 and $\lambda_{2}$= 0.8. The node and view dimensions are identical: $d_{v}=d_{n}=d=768$. We minimized changes to the baseline (DUET) [6] settings and used no additional annotations. All experiments were conducted on a single NVIDIA RTX 4090 GPU. Detailed experimental setup is provided in the supplementary material.
@@ 4.2 Comparisons with State of the Art
We conduct a comprehensive comparison on the REVERIE [41], R2R [2], and SOON [64] datasets. For fairness, we exclude methods that involve pre-exploration [66, 46] or utilize large-scale data augmentation [7, 55]. REVERIE. Table 1 presents a comparative analysis on the REVERIE. On the test unseen split, SUSA achieves an SPL score of 41.54%, significantly outperforming the previous state-of-the-art (SOTA) method by a large margin of 2.78%. Furthermore, we establish new SOTA results for RGS and RGSPL on both the validation and test unseen splits, demonstrating that the introduced textual semantics substantially elevate overall navigation, particularly benefiting object localization performance. R2R. The compared experiment results on R2R, summarized in  Tab. 2. Compared with the SOTA ESCENE [62] model on the validation and test unseen splits, SUSA achieves 64.85% and 63.83% on SPL, respectively. These gains manifest not only an improved navigation performance but also enhanced navigation efficiency, since SPL considers both success rate and trajectory length.
@@
SOON. On the SOON dataset, as shown in Tab. 3, SUSA achieves state-of-the-art performance across multiple metrics, including SR, SPL, and RGSPL, attaining SPL scores of 30.82% and 25.47% on the validation and test unseen splits, respectively. Moreover, SUSA consistently enhances the baseline (DUET)’s navigation performance and generalization on three datasets, underscoring the effectiveness of integrating textual semantics and depth exploration maps.
@@ 4.3 Diagnostic Experiment
ViT $w/o$ 51.07 46.98 33.73 32.15 23.03
CLIP $w/o$ 57.17 51.55 35.75 35.19 24.40
CLIP ResNet-50(ImageNet) 54.42 50.18 34.63 34.39 23.66
CLIP ResNet-50(Gibson) 56.72 52.00 39.21 35.05 26.54
Table 4: Ablation of different visual representations on the REVERIE validation unseen split. Gibson and ImageNet serve as pre-training sources for the depth encoder in the DSP module.
@@
1) Image Features vs. Structural Features. We analyze RGB and depth features contributions within the DSP module (Tab. 4). For RGB images, substituting ViT [11] features with those from CLIP [44] yields marked improvements across all metrics, highlighting that CLIP achieves a better understanding of the environment than its ViT-based baseline counterpart. As for depth features, we leveraged depth features extracted by ResNet-50 [17] pre-trained on the ImageNet [9] or Gibson [57] datasets to investigated the effect of depth exploration maps. Compared to ImageNet, the Gibson simulator provides spatial structural information that serves as valuable navigation priors. Therefore, in  Tab. 4, the ResNet-50 pre-trained on the Gibson extracts structural features that refine spatial perception, yielding notable improvements in the SPL metric without compromising SR or RGS. These improvements indicate that better spatial perception enhances navigation efficiency.
@@
2) Static Matching vs. Dynamic Matching. To assess the impact of static matching ($\delta=0$) and dynamic matching ($\delta=1$) strategies in the TSU module, we perform ablation studies based on balance factor $\delta$ on the R2R and REVERIE.  Tab. 5 summarizes these results, where w/o indicates the exclusion of the TSU module, corresponding to the last row results in Tab. 4. Furthermore, adaptive denotes that $\delta$ is a learnable parameter which adaptively adjusts the contribution of each matching strategy. Compared to the first row in Tab. 5, overall, incorporating the TSU module (regardless of static or dynamic matching strategy) improves SR and RGS, albeit at the cost of SPL on REVERIE. In contrast, it enhances SPL on R2R. While the TSU module can achieve superior performance under the adaptive setting, its performance is inferior when $\delta$ = 0.5. Specifically, at $\delta$ = 0.5, the SPL metric reaches 39.50% and 64.85% on the two datasets, with the RGS, nDTW, and sDTW metrics also achieving optimal performance. Excellent results demonstrate dual merits of proposed TSU module, which boosts object identification accuracy on the REVERIE task while simultaneously improving the fidelity between navigation trajectories and instructions on the R2R.
@@
Figure 5: a) and b) show the performance gap ($\downarrow$) between seen and unseen environments, while c) and d) present key metrics ($\uparrow$) for different pretraining strategies on the R2R and REVERIE.
@@
3) Improved Generalization via Pretraining. As aforementioned in Sec. 3.5, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which may negatively impact the generalization in unseen environments, particularly considering the modality differences between textual panorama, depth exploration maps, and RGB representations. As shown in Fig. 5, we observe that the partial pretraining strategy not only consistently narrow the performance gap between seen and unseen environments but also outperform complete/random pretraining strategies in navigation performance. For example, as depicted in Fig. 5 (a), the partial pretraining model achieves an SR gap of only 21.11% between the seen and unseen splits on REVERIE, reducing the gap by 5.92% compared to the complete pretraining strategy. Additionally, as shown in Fig. 5 (c), multiple metrics indicate that partial pretraining results in superior performance. Fig. 5(b) and (d) illustrate similar trends are observed on R2R, further emphasizing the generalization of the partial pretraining strategy.
ID DSP TSU HRA SR SPL RGSPL
1 46.98||||52.51 33.73||||36.06 23.03||||22.06
2 52.00||||52.53 39.21||||39.53 26.54||||24.97
3 53.17||||53.89 38.31||||40.89 26.73||||26.43
4 55.07||||51.97 39.50||||37.87 26.70||||25.20
5 51.75||||54.39 38.86||||41.54 26.56||||27.31
Table 6: Ablation study on different components of SUSA on the REVERIE (validation||||test) unseen split.
4) Overall Design. To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the REVERIE validation and test unseen split in the Tab. 6. As shown in Tab. 6, an intriguing phenomenon is observed: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates. Conversely, contrastive learning within the HRA module (#5) produces opposite outcomes, particularly with respect to the SR metric. This may be attributed to the learnable token $v$, which, while pooling hybrid environmental representations, also introduces slight noise that causes fluctuations of navigation performance. Nevertheless, the overall performance is progressively improved with the inclusion of DSP, TSU and HRA modules. Analysis of the HRA module and the computational efficiency of the entire model is presented in supplementary materials.
@@ 4.4 Qualitative Analysis
Figure 6: Predicted trajectories of SUSA and DUET on the R2R validation unseen split.
@@ 5 Conclusion
We present the SUSA architecture, which exploits complementary semantic-spatial representations beyond RGB to facilitate environmental understanding and instruction grounding. Specifically, we propose the TSU module, enabling the agent to identify the most instruction-relevant semantic. The DSP module enhances spatial awareness by grounding depth exploration maps. We further employ contrastive learning to precisely align hybrid environmental representations with instructions. Experimental results on the R2R, REVERIE, and SOON benchmarks demonstrate the effectiveness and generalization of our approach. Future work will enrich informative environmental representations to advance VLN tasks in continuous environments.
@@ 6 Detailed Experimental Setups
We evaluated the experimental results of the unseen test split on the public online leaderboards [^1] https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate. We assure the readers that we will make our code and model checkpoints publicly available.
@@ 7 Additional Experiments
Validation Unseen
Methods TL SR SPL RGS RGSPL
BEVBert [1] - 51.78 36.37 34.71 24.44
GridMM [56] 23.20 51.37 36.47 34.57 24.56
SEAT[53] - 49.45 35,51 32.83 23.14
SUSA(DP) 24.43 52.54 37.59 35.47 25.27
Depth SUSA(DEM) 22.88 52.00 39.21 35.05 26.54
KESU [15] 22.93 50.36 35.15 35.02 24.17
KERM [31] 21.85 50.44 35.38 34.51 24.45
AACL [33] 22.11 49.42 33.54 33.31 22.49
CONSOLE [34] - 50.07 34.40 34.05 23.33
Text SUSA(TP) 23.40 53.17 38.31 37.18 26.73
Table 9: Comparison with related works on the REVERIE validation unseen split.